OpenStudy (anonymous):

The sum of x and y is -16. Find x and y so that xy is a maximum?

OpenStudy (anonymous):

the answers i have are : (0, -16) , (-8,-8), (8,8) or (-16,0)

OpenStudy (anonymous):

im not in calculuss. im in algebra 2.

Parth (parthkohli):

Why Calculus? Why not just try out the answer choices?

OpenStudy (anonymous):

Let the 2 numbers be x and y. so, we are told that x+y=-16 --eqn1 and that xy is to be a maximum. so we have the function f(x) = xy. Now get rid of y by using eqn1 (written as y = -16-x... so f(x) = x(-16-x) We are told to find those 2 numbers that have the largest product dy/dx = -16 - 2x = 0 so, x=-8 so therefore, y must be -8 too this is the answer: x=y=-8

OpenStudy (anonymous):

(-8,-8)

Parth (parthkohli):

That is making it too complicated.

OpenStudy (anonymous):

@noniebug u got it

Parth (parthkohli):

(8,8) is not the answer, as it sums up to 16.

Parth (parthkohli):

Now see what these are: -8 x -8 -16 x 0 0 x -16

OpenStudy (anonymous):

so its (-8,-8) ?

OpenStudy (anonymous):

easy to plug the values and see as prath told

OpenStudy (anonymous):

if the options are not given go by my method!

OpenStudy (anonymous):

i think this is Compatible with algebra for all real numbers a,b \( (a+b)^2 \ge 4ab \) and equality occurs when \(a=b\) proof: we know that \((a-b)^2 \ge 0 \) so \(a^2+b^2 \ge 2ab \) and \(a^2+b^2+2ab \ge 4ab \) finally we have \( (a+b)^2 \ge 4ab \). equality occurs when \((a-b)^2 = 0 \) or \(a=b \). on the other hand \( (a+b)^2 \ge 4ab \) tells us \( max[4ab]=(a+b)^2 \) and it happens when \(a=b\). now for ur case \(xy\) is maximum so \(x=y\) and \(2x=-16\) finally \(x=y=-8 \)